OpenStudy (anonymous):

Limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow ∞} \frac{ x^2e^x }{ x^5+1 }\]

OpenStudy (anonymous):

Limit does not exist

OpenStudy (ikram002p):

its exist shail

OpenStudy (vishweshshrimali5):

Yes the limit exists it is \(\infty\) , right ? @ikram002p

OpenStudy (vishweshshrimali5):

I am not so sure, did you get the same ?

OpenStudy (anonymous):

\[\lim_{x \rightarrow ∞}\frac{ e^x }{ x^3+\frac{ 1 }{ x^2 }}=L'H=\frac{ e^x }{ 3x^2-\frac{ 2 }{ x^3 } }=L'H=\frac{ e^x }{ 6x+\frac{ 6 }{ x^4 } }=L'H=\frac{ e^x }{ 6-\frac{ 24 }{ x^5 } }=+∞\] Thank you guys

OpenStudy (vishweshshrimali5):

Well yeah that's about what I did @Alexander95

OpenStudy (ikram002p):

xD sry guys i thought x goes to 0 my bad

OpenStudy (vishweshshrimali5):

Np